Acoma, Inc., has determined a standard direct materials cost per unit of $8 (2 feet times $4 per foot). Last month, Acoma purchased and used 4, 200 feet of direct materials for which it paid $15, 750. The company produced and sold 2,000 units during the month.
 
Calculate the direct materials price, quantity, and spending variances.
Answer #1
Direct material price variance = Actual cost-(Standard rate*Actual Quantity)
=15750-(4*4200)= 1050 Favorable
Direct material quantiy variance = Standard rate*(Standard quantity-Actual Quantity)
=4*(4000-4200)= 800 Unfavorable
Direct material spending variance = 1050F + 800U = 250F